### v1

# Hugo Malard1  Michel Olvera1  Stéphane Lathuiliere1  Slim Essid1 1LTCI, Télécom Paris, Institut Polytechnique de Paris

Abstract
Large-scale pre-trained audio and image models demonstrate an unprecedented degree of generalization, making them suitable for a wide range of applications. Here, we tackle the specific task of sound-prompted segmentation, aiming to segment image regions corresponding to objects heard in an audio signal. Most existing approaches tackle this problem by fine-tuning pre-trained models or by training additional modules specifically for the task. We adopt a different strategy: we introduce a training-free approach that leverages Non-negative Matrix Factorization (NMF) to co-factorize audio and visual features from pre-trained models to reveal shared interpretable concepts. These concepts are passed to an open-vocabulary segmentation model for precise segmentation maps. By using frozen pre-trained models, our method achieves high generalization and establishes state-of-the-art performance in unsupervised sound-prompted segmentation, significantly surpassing previous unsupervised methods.
1 Introduction
Audio-visual perception has attracted considerable interest due to its practical applications in various fields, especially robotics, video understanding [4, 39], and recently audio-visual language modeling [25, 6]. A particularly compelling task in this area is sound-prompted segmentation, which consists in pinpointing the image regions associated with sounds in a corresponding audio signal. Most approaches proceed to align audio and visual representations through contrastive learning techniques [41, 38, 30]. These techniques leverage paired audio-visual data to learn a shared embedding space, where corresponding audio and visual features are brought closer together, while non-corresponding pairs are pushed apart. The main challenge lies in learning local features that can be used for segmenting audio-related concepts, as videos typically provide global-level alignment (as opposed to local alignment where each part of the image is matched with a part of the audio). More recently, a few works have made significant progress in this direction through the use of pre-trained models. For example, Park et al. [31] use a pre-trained CLIP encoder [32] and audio tokenizer within a relatively complex framework that translates audio signals into tokens compatible with CLIP’s text encoder. This is combined with a grounding mechanism, enabling audio-driven segmentation of the image. Similarly, Hamilton et al.’s DenseAV  [17] identifies image regions linked to sounds without supervision. Using a DINO image backbone [1] and the HUBERT audio transformer [19], it employs multi-head feature aggregation for contrastive learning on video-audio pairs.
Figure 1: Our method takes a representation of an image and its associated audio as input, decomposing them into a product of ’semantic’ factors and (spatial or temporal) activations. This decomposition enables locating parts of the original input corresponding to the concept present in both the image and the audio.
While these approaches tackle sound-prompted segmentation through carefully designed architectures and dedicated training procedures, we propose a radically different strategy. Specifically, we ask: can this task be addressed using only pre-trained audio and image models, without any task-specific training, by directly analyzing their feature representations?
To answer this question, we introduce TACO: a Training-free Audio-visual CO-factorization approach, based on soft Non-negative Matrix co-Factorization (soft co-NMF) which identifies audio and visual tokens that are co-activated in the input signal. By leveraging pre-trained CLIP and CLAP [10] backbones for image and audio, our co-NMF formulation enables audio-visual correspondence analysis without the need for training. Specifically, using a set of semantic anchor words projected into the CLIP and CLAP spaces, we are able to match audio features with visual ones in an interpretable fashion. Through simple inference-time optimization, our framework segments both visual and audio signals, performing object segmentation and sound localization in a training-free manner (illustrated in Figure 1). Furthermore, the interpretable factors identified by our decomposition framework allow us to directly prompt a pre-trained open-vocabulary segmenter, such as FC-CLIP [47], enhancing segmentation quality while remaining zero-shot.
Our approach stands out for its training-free paradigm and the interpretability offered by the NMF framework. Additionally, leveraging the robustness of the frozen CLIP and CLAP models enables us to achieve significantly superior performance compared to existing unsupervised training-based methods on established benchmarks. To summarize, our contributions are:
  • We introduce TACO, the first training-free approach for sound-prompted segmentation based on pre-trained deep representations. Our framework leverages the inherent interpretability of NMF, enabling clear, semantically grounded visualization of interactions between segmentation outputs and the concepts identified within the signal.
  • We enhance the quality of the segmentation provided by our co-NMF framework by incorporating a pre-trained open-vocabulary model (namely, FC-CLIP [47]), prompting its decoder with the most activated factors identified by our NMF system, for improved accuracy.
  • We rigorously validate our approach through extensive experiments, providing both quantitative and qualitative analyses. We demonstrate the superior performance of our training-free approach across four datasets and multiple variants of the sound-prompted segmentation task. Additionally, the code used in our experiments will be released.
2 Related Work
2.1 Sound-Prompted Segmentation
Sound-prompted segmentation is the task of localizing in an image the object emitting the sound that is heard on an associated audio. Recent methods primarily utilize cross-modal attention for sound-prompted segmentation [42, 36, 35], often combined with a contrastive loss that learns to align audio and image global representations, hypothesizing that this process would implicitly foster the emergence of local alignment. Extending this contrastive learning framework, numerous enhancements have been introduced. These include integrating challenging negatives from background areas [3], adopting iterative contrastive learning with pseudo-labels from earlier model epochs [22], maintaining transformation invariance and equivariance through geometric consistency [23], leveraging semantically similar hard positives [37], and implementing false negative-aware contrastive learning through intra-modal similarities [40].
Very recently, large pre-trained models have started to be used in order to benefit from their rich data representation. ACL-SSL [31] relies on the CLIP image encoder as well as a segmentation model CLIP-Seg[24], and learns an additional module that projects the audio in the CLIP space. Relying on DINO and HUBERT (further fine-tuned), Hamilton et al.[17] trained a multi-head module that identifies the localization of the sound in the image features.
Unlike these methods, our framework relies on a training-free setup leveraging pre-trained models and matrix decomposition. By avoiding training, our method not only avoids the need for a training dataset but also preserves the generalization abilities of the pre-trained model.
2.2 NMF and Deep Feature Factorization
Non-negative matrix factorization (NMF) has found applications in diverse fields, including audio source separation [15], document clustering [44], or face recognition [16], to mention a few. Previous research has expanded NMF to multiple layers [7], implemented NMF using neural networks [9], and utilized NMF approximations as inputs to neural networks[43]. More recently it has been applied to decompose deep neural features, for concept discovery [8], interpretability [27, 28] or audio-visual source separation [29]. While it has already been applied to co-factorize audio and visual (handcrafted) features [34], to the best of our knowledge this work is the first one to apply it to deep features for sound-prompted segmentation, as well as the first attempt to co-factorize representations from different spaces while enforcing semantic matching.
3 [Uncaptioned image] TACO: Training-free Audio-visual CO-factorization
This section introduces our method, TACO, which allows training-free sound-prompted segmentation. Section 3.1 starts by presenting the main motivations. Section 3.2 provides the necessary background on NMF, and section 3.3 introduces our framework for extracting audio-visual concepts using NMF, while section 3.4 describes our complete solution for audio-source segmentation, which integrates our factorization approach into a pre-trained open-vocabulary image segmenter, enabling fine segmentation in a training-free manner.
3.1 Motivation
Our methodology leverages the capacity of recent large pre-trained models to effectively localize patterns within their input signals [47]. Localization here refers to identifying spatial regions in images and temporal segments in audio. From this observation, we hypothesize that separately trained audio and visual encoders contain all the necessary information for sound-prompted segmentation without requiring additional training. To achieve this, we propose a method, TACO, which identifies concepts that co-activate across audio and image representations exploiting the Non-negative Matrix Factorization (NMF) framework [21]. NMF provides a training-free approach that jointly decomposes each signal into an activation matrix and a concept matrix. The activation matrix, in particular, can be interpreted as a segmentation matrix, highlighting regions of concept alignment.
Before detailing our approach, it is important to note that NMF is a tool that is meant to interpret input matrices composed of only non-negative values. The non-negativity constraints are key to the interpretability offered by the NMF framework. However, deep representations often include negative coefficients. To address this, a straightforward solution is to clip all negative values to zero, thus ensuring matrix non-negativity. Although this adjustment may result in some information loss, our preliminary experiments indicate that with backbones like CLIP [32] and CLAP [10], the representation power remains unaffected for our specific tasks and settings. Further experimental details on this are provided in Sec. 4.
3.2 Background
NMF
Non-negative Matrix Factorization (NMF) [21] involves decomposing a non-negative matrix $X\in\mathbb{R}^{N\times C}_{+}$ into two non-negative matrices $U\in\mathbb{R}^{N\times K}_{+}$ and $V\in\mathbb{R}^{K\times C}_{+}$ such that $X\approx UV$. In typical machine learning applications of NMF [12], $N$ represents the number of observations and $C$ denotes the number of (non-negative) features per observation, while $K$ is the rank of the factorization, controlling the number of factors or components into which the original matrix is decomposed. Assuming a measure of fit $D$ is considered, NMF can be formulated as the following constrained optimization problem:
$\min_{U,V\geq 0}D(X|UV).$ (1)
As the $V$ matrix is interpreted as the factor matrix, $U$ corresponds to the activations of those factors.
Co-NMF
extends the NMF framework to multiple data views analyzed in parallel. Assuming two data matrices $X_{1}\in\mathbb{R}^{N\times C}_{+}$ and $X_{2}\in\mathbb{R}^{N\times C}_{+}$ containing different modalities (such as audio and visual streams of a video), co-NMF exploits the mutual information shared between these modalities. It assumes that the different modalities share a common activation matrix $U$. In particular, [45, 46] decomposes these matrices into a shared activation matrix $U\in\mathbb{R}^{K\times C}_{+}$ with view-specific factors $V_{1}\in\mathbb{R}^{N\times K}_{+}$ and $V_{2}\in\mathbb{R}^{N\times K}_{+}$. This formulation forces the activation matrix $U$ to be consistent across both modalities, reflecting the intuition that events occurring in the audio also happen in the same time in the image stream. Although the factors are encoded in different spaces, they should remain the same. Formally, assuming measures $D_{1}$ and $D_{2}$, and weighting hyperparameters $\beta_{1}>0$, $\beta_{2}>0$, co-NMF solves the following problem:
$\min_{U,V_{1},V_{2}}\beta_{1}D_{1}(X_{1}|UV_{1})+\beta_{2}D_{2}(X_{2}|UV_{2}),$ (2)
subject to non-negativity constraints on $U$, $V_{1}$, and $V_{2}$.
Note that, as there might be some temporal shift between the audio and image events, enforcing the same $U$ matrix for both decompositions can be overly restrictive. To mitigate this issue, soft non-negative Matrix Co-Factorization (soft co-NMF) [34] introduces two separate activation matrices and includes a penalty term to minimize their dissimilarity. It is obtained by solving the following optimization problem, with a penalty function $P$, and hyperparameter $\beta_{p}>0$
$\min_{U_{1},U_{1},V_{2},V_{2}}\beta_{1}D_{1}(X_{1}|U_{1}V_{1})+\beta_{2}D_{2}(X_{2}|U_{2}V_{2})$ (3)
$+\beta_{p}P(U_{1},V_{1},U_{2},V_{2}),$
subject to non-negativity constraints on $U_{1}$, $U_{2}$, $V_{1}$ and $V_{2}$.
3.3 Deep Non-negative Matrix co-Factorization
This section details how TACO leverages soft co-NMF to perform a zero-shot sound-prompted segmentation.
Figure 2: Complete pipeline: both the audio and the image are encoded using their respective encoder and their representations are used to perform the co-NMF. FC-CLIP is prompted using the image factors ($V_{I}$) and the segmentation corresponding to the sounding image factor ($V_{I}^{k\star}$) is kept as the final segmentation.
Soft co-NMF for token decomposition
We assume access to an input video with a corresponding audio signal. For simplicity, let us initially consider the central frame of the video; we will later extend our approach to handle the entire video, incorporating the temporal dimension. Additionally, we assume the availability of trained audio and image backbones. Feeding the audio signal into the audio backbone produces a matrix $X_{A}\in\mathbb{R}^{N_{T}\times C_{A}}_{+}$, where $N_{T}$ is the number of tokens, and $C_{A}$ is the token dimension. Similarly, the image encoder maps an image to a feature representation $X_{I}\in\mathbb{R}^{HW\times C_{I}}_{+}$, where $H$ and $W$ represent the spatial dimensions, and $C_{I}$ is the number of channels. Note that the spatial dimensions are flattened, resulting in a 2D matrix compatible with NMF inputs. Note that our approach diverges from prior works as the image activations encapsulate spatial, rather than temporal, information.
By applying soft co-NMF, we aim at factorizing $X_{A}$ into matrices $U_{A}\in\mathbb{R}^{N_{T}\times K}_{+}$ and $V_{A}\in\mathbb{R}^{K\times C_{A}}_{+}$, along with $X_{I}$ into $U_{I}\in\mathbb{R}^{HW\times K}_{+}$ and $V_{I}\in\mathbb{R}^{K\times C_{I}}_{+}$, where $K$ denotes the predefined number of factors.
To enforce the non-negativity constraints in Eq. (3), we re-parameterize the problem using the sigmoid function $\sigma$ and optimize matrices $\tilde{U}_{A}\in\mathbb{R}^{N_{T}\times K}$ and $\tilde{U}_{I}\in\mathbb{R}^{HW\times K}$, such that $U_{A}=\sigma(\tilde{U}_{A})$ and $U_{I}=\sigma(\tilde{U}_{I})$ with $U_{I}\in[0,1]^{HW\times K}$ and $U_{A}\in[0,1]^{N_{T}\times K}$. Restricting the decomposition to the interval $[0,1]$ rather than $\mathbb{R}_{+}$ offers the advantage of an interpretable, well-calibrated scale. Specifically, the $k$-th column of $U_{I}$ and $U_{A}$ can be interpreted as the activation of the $k^{th}$ factor for each token. In this way, the Hadamard product between the original tokens and a column of $U_{I}$ or $U_{A}$ can be viewed as soft masking, highlighting the portion of the original input that activates the $k^{th}$ factor. Overall, those shifts lead us to challenge the existing soft co-NMF framework, proposing instead that audio and image factors should align, while activations may differ. Thus, we define a specific penalty function to impose semantic closeness between the factor matrices. As the modality-specific factors lie in different spaces, we introduce a new manner to ensure their semantic closeness.
Penalty function
Applying soft co-NMF to a specific problem requires careful definition of the penalty term $P$ in Eq. (3), as it impacts both the quality of the factorization and its interpretability. To achieve a semantically consistent factorization of audio-visual observations, $P$ must effectively capture the semantic similarity between $V_{I}$ and $V_{A}$. However, since the backbones (e.g., CLIP and CLAP) are not aligned, audio and image factors are encoded in two different subspaces, making $L_{1}$ or $L_{2}$ norms unsuitable as penalty functions. Moreover, audio and image representations are not necessarily commensurable: they often differ in dimensionality, precluding direct use of standard distances.
To address this, we propose projecting the audio and visual feature representations into a unified feature space to enable direct comparison. Specifically, we estimate aligned semantic descriptors $\mathcal{D}_{I^{k}}$ and $\mathcal{D}_{A^{k}}$ for the $k^{th}$ column of $U_{I}$ and $U_{A}$, respectively. These descriptors are computed using $J$ semantic anchors, $\{b_{I}^{j}\}_{j=1}^{J}$ and $\{b_{A}^{j}\}_{j=1}^{J}$, which lie in the same spaces as $X_{I}$ and $X_{A}$. Each anchor pair $(b_{I}^{j},b_{A}^{j})$ represents a shared semantic concept, enabling the estimation of aligned semantic descriptors.
Specifically, to estimate aligned semantic anchors, we utilize CLIP [10] and CLAP [32] as image and audio backbones. Since these encoders are aligned with text encoders, we can exploit their respective text embeddings to align identical words across audio and image spaces. Thus, we define $b_{I}^{j}=\mathcal{E}_{I}(t_{j})$ and $b_{A}^{j}=\mathcal{E}_{A}(t_{j})$, where $\mathcal{E}_{I}$ and $\mathcal{E}_{A}$ are the text encoders of CLIP and CLAP, respectively, and $\{t_{j}\}_{j=1}^{J}$ represents a predefined word bank.
Figure 3: CLIP and CLAP encode audio and images in different spaces, complicating distance computation. Our method makes use of semantic anchors to project semantic components (pre-trained representations soft masked by the activations $U_{I}^{k}$ and $U_{A}^{k}$ obtained during the co-factorization) in an audio-visual semantic space where standard distances can be applied. Those distances are used as a penalty function during the decomposition to ensure that the audio and image concepts are the same.
We now describe the process of estimating the semantic descriptors from the matrices $U_{A}$ and $U_{I}$ using the aligned semantic anchors. For simplicity, we detail only the estimation of semantic image descriptors; analogous steps apply to the audio descriptors.
First, we introduce semantic components $\mathcal{C}_{I}^{k}$ associated to each column $k$ of $U_{I}$. Since $U_{I}^{k}$ represents the spatial activation of the $k^{th}$ factor (corresponding to $V_{I}^{k}$), it can be interpreted as a soft segmentation mask for the input $X_{I}$. Applying these soft-masks to $X_{I}$ and averaging the resulting representations extracts the concept associated with the $k^{th}$ factor from the image features. Formally, the $k^{th}$ image semantic component is computed as follows:
$\mathcal{C}_{I}^{k}=\text{avg}(X_{I}\odot U_{I}^{k}),$ (4)
[…]
$\mathcal{D}_{I^{k}}=\begin{pmatrix}\cos(\mathcal{C}_{I}^{k},b_{I}^{1})\\ \vdots\\ \cos(\mathcal{C}_{I}^{k},b_{I}^{J})\end{pmatrix},$ (5)
where $\cos(.)$ denotes cosine similarity. Intuitively, a descriptor quantifies the alignment of the semantic component with each semantic anchor (illustrated in Figure 3).
Note that $V_{I}^{k}$ could have been used to play the role of semantic components instead of $\mathcal{C}_{I}^{k}$. However, $V_{I}^{k}$ is less resilient to variations of the semantic anchors. Indeed, during the NMF optimization, $V_{I}^{k}$ can take any range of values, potentially degenerating to values similar to those of $b_{I}$, which would be equivalent to performing a simple classification. In contrast, each $\mathcal{C}_{I}^{k}$ is inherently constrained to be a weighted average of elements from the original representations, thus discouraging it from merely replicating values from $b_{I}$. We ablate this choice in Section 4.
Audio semantic descriptors $\mathcal{D}_{A^{k}}$ are computed in the same way as their image counterparts. The penalty function is then defined as the cross-entropy between the image and audio semantic descriptors. However, enforcing similarity across all semantic descriptors may be detrimental, as certain semantic components may be visually present but inaudible, or vice versa. To address this, alignment is applied only to the closest audio and image semantic components. This choice is ablated in Section 4. Using the $L_{2}$ norm for reconstruction, we obtain $V_{A}$, $V_{I}$, $U_{A}$, and $U_{I}$ by solving the following optimization problem:
$\min_{V_{A},V_{I},U_{A},U_{I}\geq 0}\big{[}\beta_{1}\|X_{A}-U_{A}V_{A}\|_{2}^{2}$ (6)
$+\beta_{2}\|X_{I}-U_{I}V_{I}\|_{2}^{2}$
$+\beta_{p}\min_{k}\,\text{CE}(\mathcal{D}_{I^{k}},\mathcal{D}_{A^{k}})\big{]},$
where CE denotes the cross-entropy function.
Interpretation
After optimization, the $k^{\text{th}}$ row of $U_{I}$ represents the segmentation of the $k^{\text{th}}$ factor in $V_{I}$ (i.e. its $k^{\text{th}}$ column) within the original image. The index $k$ that realizes $\min_{k}\,\text{CE}(\mathcal{D}_{I^{k}},\mathcal{D}_{A^{k}})$ corresponds to the dominant semantic component shared between the audio and image modalities, denoted as $k^{\star}$. Consequently, row $k^{\star}$ of $U_{I}$, referred to as $U_{I}^{k^{\star}}$, can be interpreted as the segmentation of the image region associated with the sound in the audio signal. Likewise, the segmentation of the spectrogram for this same semantic component is represented by $U_{A}^{k^{\star}}$.
Temporal consistency
To incorporate the temporal dimension in both the audio and visual inputs, if dealing with a video sequence, we introduce a temporal consistency regularization term in the decomposition process. Given a video, we extract $T$ frames at a constant interval, and construct $T$ audio-image pairs, denoted by $\{X_{A_{t}}\}_{t=1}^{T}$ and $\{X_{I_{t}}\}_{t=1}^{T}$, by extracting the $T$ consecutive non-overlapped audio frames centered around an image frame position. For each frame pair, we perform a decomposition into matrices $V_{A_{t}},V_{I_{t}},U_{A_{t}},U_{I_{t}}$, introducing a regularization term $\mathcal{R}_{t}$ to encourage consistency between the primary shared factors, $V_{I_{t}}^{k^{\star}}$ and $V_{A_{t}}^{k^{\star}}$, across consecutive frames. Formally, we add the following $\mathcal{R}_{t}$ regularization term to the optimization objective of Eq. (6):
$\mathcal{R}_{t}=-\beta_{t}\sum_{t=1}^{T}\cos(V_{I_{t}}^{k^{\star}},V_{I_{t+1}}^{k^{\star}})+\cos(V_{A_{t}}^{k^{\star}},V_{A_{t+1}}^{k^{\star}}),$ (7)
where $\beta_{t}\geq 0$ controls the temporal regularization.
3.4 Improved Segmentation with an Open Vocabulary Segmenter
Our soft co-NMF approach generates a segmentation mask for objects producing sound in the audio input through the $k^{\star}$-th row of $U_{I}$. However, the segmentation accuracy is limited, as the CLIP features in the visual input $X_{I}$ were trained only for image-level semantic alignment, not for precise segmentation. To enhance segmentation quality, we propose to integrate a pre-trained open-vocabulary segmentation model. To maintain compatibility with our NMF framework and keep the approach zero-shot, we select a model based on CLIP that preserves alignment with the text encoder without additional fine-tuning of the CLIP encoder. Here, we use the FC-CLIP model [47], which segments images using CLIP’s image encoder via text prompts encoded using the CLIP text encoder.
A valuable feature of FC-CLIP is its shared CLIP embedding space for both image and text inputs, which allows it to be prompted by vectors from the image space. Leveraging this, we prompt the FC-CLIP decoder with factors identified by our NMF framework. More precisely, in our soft co-NMF decomposition, the rows of matrix $V_{I}$ represent the factors in the image space. Therefore, we can use them to prompt FC-CLIP, treating each factor in $V_{I}$ as a segmentation class. Specifically, $V_{I}^{k^{\star}}$ describes the shared “sounding” factor between audio and visual inputs.
Since $V_{I}^{k^{\star}}$ resides in the CLIP space, it can also be compared to class name embeddings, allowing us to infer the segmented object’s class names and achieve sound prompted image semantic segmentation. The complete pipeline is depicted in Figure 2. Note that, incorporating FC-CLIP is computationally efficient, as the CLIP representation is computed only once before the soft co-NMF decomposition, with only the lightweight FC-CLIP decoder adding minor additional processing overhead (there is $\sim$200M parameters for the CLIP encoder and $\sim$20M for the FC-CLIP decoder).
4 Experimental protocol
Implementation details. To extract the modality-specific representations we utilize the MSCLAP 2023 [10] ($\sim$80M parameters) audio backbone and the OpenCLIP [5] ConvNeXt-Large CLIP trained on LAION 2B [33] ($\sim$200M parameters). For all experiments, we fix the decomposition coefficients to $\beta_{I}=\beta_{2}=1,\beta_{p}=125$, and $K=8$. For the temporal consistency constraint, we set the image and audio frame rate to one and use $\beta_{t}=1$. However, since temporal consistency is specifically designed for single-source segmentation, we set $\beta_{t}=0$ for the multi-source task. Those hyperparameters were selected based on preliminary experiments conducted on the AVSBench validation set. The optimization problem is solved using gradient descent in 1800 steps with a learning rate of 0.25. Importantly, the optimization process induces minimal computational overhead compared to the use of all pre-trained modules (details provided in the supplementary materials).
Datasets. We evaluate our method on five datasets: three for sound-prompted segmentation and two for sound-prompted semantic segmentation. For sound-prompted segmentation, we use: AVS-Bench [49], with binary segmentation maps for audio-visually related pixels, divided into single-source (S4,  5K samples) and multi-source (MS3,  400 samples) datasets. Additionally, we use ADE20k sound-prompted [17], which pairs ADE20k segmentation masks and images[48] with audio from VGGSound [2]. For sound-prompted semantic segmentation, we created ADE Sound Prompted Semantic: a semantic version of the ADE20k sound-prompted dataset, by assigning class labels to segmentation masks (detailed in supp. material). Moreover, we conduct experiments on AVSS [50], ($\sim$11K samples) comprising single and multi-source sounding objects.
Evaluation metrics. We adopt standard segmentation metrics for evaluation: mean Intersection over Union (mIoU) and F-score for the semantic segmentation tasks, and mean Average Precision (mAP) as well as mask Intersection over Union (mask-IoU) and F-score for the binary segmentation tasks (we distinguish between mask and mean IoU to ensure comparability with previous works). Each experiment involving NMF decomposition was conducted three times, with the results reported as the average and standard deviation of the scores.
5 Results and Discussion
In this section, we compare our approach with other unsupervised methods tackling the task of sound-prompted segmentation. Most of the methods rely on fine-tuning (even if done unsupervisedly) which significantly modifies the original representations of pre-trained models. In contrast, our method is training-free and fully preserves the large generalization ability of the models it builds upon.
For semantic segmentation, we use the class names provided with the dataset as the word bank. For non-semantic segmentation, no predefined class set is needed. We consider two variants of our method: in-domain segmentation, which uses a dataset-specific class set, and out-of-domain segmentation, which relies on a general set of 527 AudioSet[11] tags. For fair comparison, we report results using the general class list, unless otherwise noted.
ADE SP S4 MS3 AVSS ADE SP S.
A-FC-CLIP 45.31 51.52 33.67 11.95 24.95
A+-FC-CLIP 41.89 51.63 31.67 11.82 23.73
Table 1: Impact of clamping on the representation before feeding FC-CLIP. We report mIoU on five datasets. ADE SP S. corresponds to the ADE SP Semantic segmentation dataset. mask-IoU is reported for S4 and S3 and mIoU for all the other datasets.
Baseline. In the absence of prior work on unsupervised sound-prompted semantic segmentation, and to have a fair competitor that relies on the same pre-trained model, we consider a straightforward yet effective baseline approach: Audio-prompted-FC-CLIP (A-FC-CLIP). This method takes an audio input and its corresponding image, utilizing CLAP to determine the most likely class from the dataset’s class pool. FC-CLIP is then prompted to segment this class within the image. Note that for the baseline we use the dataset-specific class list as we use CLAP as a classifier that should rely on predefined class names.
Method S4 MS3
Training-free mask-IoU $\uparrow$ F-Score $\uparrow$ mask-IoU $\uparrow$ F-Score $\uparrow$
SLAVC [26] 28.10 34.60 24.37 25.56
MarginNCE [30] 33.27 45.33 27.31 31.56
FNAC [38] 27.15 31.40 21.98 22.50
Alignment [41] 29.60 35.90 - -
ACL-SSL [31] 59.76 69.03 41.08 46.67
A-FC-CLIP 51.52 57.25 33.67 35.87
TACO 64.04$\pm$ 0.25 71.50$\pm$0.20 50.46 $\pm$3.12 58.11$\pm$2.95
Table 2: Quantitative results of sound-prompted segmentation on the AVS-Bench test sets.
Preliminary experiments: impact of clamping. To test the resilience of CLIP representations to clamping negative values, we experiment with FC-CLIP prompted only with the nonnegative part (via ReLU) of the CLIP text encoding, which is referred to as A+-FC-CLIP. Table 1 shows the performance degradation across all datasets. Except for the ADE SP dataset ($\sim$3.5 point drop), the decrease never exceeds 2 points. This experiment validates the use of clamping, enabling the application of NMF-based methods so as to benefit from their interpretability.
5.1 Sound-prompted Segmentation
We now discuss experiments on (non-semantic) Sound-Prompted Segmentation. Table 2 compares our approach with other unsupervised methods on the AVS-Bench, in both single and multi-source setups. TACO demonstrates strong performance relative to unsupervised training-based alternatives, outperforming the best competitor method by $\sim$5 and $\sim$9 points of mIoU in S4 and MS3 tasks. Our method also largely outperforms the A-FC-CLIP baseline that uses the same pre-trained models, underlining the effectiveness of our decomposition approach.
In the multi-source setup, while baseline performance drops substantially due to its limitation in handling only a single concept, our method shows robust performance. We hypothesize that robustness arises because $V_{I}^{k^{\star}}$ encodes information about all the sounding objects rather than just one, within the sounding factor The hypothesis is explored qualitatively in Section 5.3.
Table 3 presents performance results on the recently proposed ADE Sound Prompted (ADE SP) dataset, along with variations of our approach on the same dataset to understand its gains. We compare existing methods with TACO variants, and our decomposition without FC-CLIP. Our experiments show that using $U_{I}^{k^{\star}}$ alone from the decomposition leads to better segmentation of the image segments compared to previous audio-visual models. This is noteworthy since all competitors (like DenseAV [17], CAVMAE [14], DAVENet [18], and ImageBind [13]) are fine-tuned for this specific task. In contrast, we use frozen features, decompose them to identify matching factors, and demonstrate that these features contain the spatial information necessary for localization. Predictably, the use of FC-CLIP segmenter substantially refines the segmentation and the scores. Now, we examine the different variants of our approach to understand the large performance gap between TACO and existing methods. In the TACO w/o component variant, using $V_{I}^{k}$ and $V_{A}^{k}$ instead of the soft mask representations as semantic components during the optimization process (i.e., using $\mathcal{D}_{I}^{k}\!=\!V_{I}^{k}$ and $\mathcal{D}_{A}^{k}\!=\!V_{A}^{k}$) leads to degraded performance. In TACO No-min, we force all the concepts to be close, instead of only matching the closest, by using a penalty term where the min operator in (6) is replaced by an average. This results in a noticeable drop in performance. The last variant, TACO w/o penalty, shows the results using $\beta_{p}\!=\!0$ (i.e., without penalization), which is equivalent to performing two independent decompositions and choosing $k^{\star}\!=\!\operatorname{argmin}_{k}CE(\mathcal{D}_{I}^{k},\mathcal{D}_{A}^{k})$. As this decomposition is uni-modal, performance significantly drops.
[…]
DAVENet [18] - 17.0 16.8
CAVMAE [14] - 20.5 26.0
ImageBind [13] - 18.3 18.1
DenseAV [17] - 25.5 32.4
$U_{I}^{k^{\star}}$ (Proposed) - 27.74$\pm$0.48 35.75$\pm$0.33
A-FC-CLIP FC-CLIP 45.31 50.36
TACO No-min FC-CLIP 29.93$\pm$1.95 34.59$\pm$3.17
w/o component 22.70$\pm$0.74 23.68$\pm$2.03
w/o penalty 26.75$\pm$0.14 30.17$\pm$0.12
Proposed 51.57$\pm$0.86 60.01$\pm$1.38
Table 3: Quantitative results of sound-prompted segmentation on the ADE SP dataset.
Word Bank S4 MS3 ADE SP
$U_{i}^{k^{\star}}$ General 27.78$\pm$0.06 28.20$\pm$0.61 32.42$\pm$0.74
Specific 28.76$\pm$0.06 24.94$\pm$0.62 27.74$\pm$0.48
TACO Specific 64.73$\pm$0.16 47.23$\pm$2.25 54.05$\pm$1.40
General 64.04$\pm$0.25 50.45$\pm$3.12 51.57$\pm$0.86
Table 4: Comparison of the two word banks on sound-prompted segmentation datasets. mask-IoU is reported for S4 and MS3, while mIoU is reported for ADE SP.
Impact of the Word bank. Table 4 presents the mask-IoU scores on the three used sound-prompted segmentation datasets using either a general class list or one specific to the dataset. Using a specific class list overall improves both TACO results (+1 point on S4 and 3.5 points on ADE SP) and the segmentation using $U_{i}^{k^{\star}}$ (+1 point on S4 and 5 points on ADE SP). On the multi-source MS3 dataset, the general class list performs better. This can be explained by the fact that, when multiple sound sources are present, a broader set of concepts can be beneficial, as one may align with the union of the sources (e.g., Family’ could encompass both a woman and a baby making sounds).
5.2 Sound-Prompted Semantic Segmentation
Sound-Prompted Semantic Segmentation differs from Sound-Prompted Segmentation seen in the previous section as it requires the estimation of class labels corresponding to the mask. In TACO, classifying the audio prompt $V_{I}^{k^{\star}}$ allows us to obtain the label of the segment. We run experiments on the AVSS and ADE20k sound-prompted semantic datasets. As shown in Table 5, our method greatly outperforms the baseline on the two datasets as it can take advantage of both the image and the audio decompositions (+8 and 15 points of mIoU in AVSS and ADE SP Semantic, respectively). On the ADE SP Semantic, which contains a single class per image, the method performs well. However, on the AVSS dataset that contains samples with multiple classes, the performances drop. We attribute this to the class decoding process: since we estimate the class label as the closest one to the value of $V_{I}^{k^{\star}}$, we consider the full segmentation as a single class. Nonetheless, the class-agnostic segmentation masks are accurate as the decomposition encodes all the concepts in a single vector (shown in supplementary materials).
Method AVSS ADE SP Semantic
mIoU $\uparrow$ F-Score $\uparrow$ mIoU $\uparrow$ F-Score $\uparrow$
A-FC-CLIP 11.95 13.56 20.91 24.95
TACO 20.50 $\pm$0.10 23.02 $\pm$0.09 37.70$\pm$1.61 44.39$\pm$1.92
Table 5: Quantitative results of sound-prompted semantic segmentation on the AVSS and ADE SP Semantic datasets.
5.3 Qualitative Analysis
Figure 4 shows images along with their associated $V_{I}^{k^{\star}}$ reshaped as a matrix, the segmentation from TACO, and an illustration of the associated audio. The first row shows a challenging example from the S4 dataset where the sound source is nearly invisible, yet both TACO and $V_{I}^{k^{\star}}$ successfully segment it. While recent works [20] underline how audio-visual segmenter models fail to deal with negative audio (i.e. audio depicting an event not present in the image), the second row illustrates the robustness of our method on that particular problem. As there is no common concept between the audio and image (given that the sound track of this example features music instead of the original audio recording of the scene), the estimated $V_{I}^{k^{\star}}$ contains only noise, and as FC-CLIP does not find the concept, hence the final segmentation is empty, showing the resilience of TACO to those negative audios. Finally, the last row illustrates the case of multiple potential sources in the image. Even though the image features a Guitar, a Djembe, and singers, the model correctly segments only the people as only singing voice is heard on this excerpt.
Figure 4: Qualitative segmentation examples from ADE SP and S4 datasets.
Analysis in the Multi-Source setting. Figure 5 shows a segmentation example from the MS3 dataset as well as the similarity of the associated $V_{I}^{k^{\star}}$ with some words. In the first row, the sound heard is a baby babbling, while the image contains both the baby, a woman, and a sofa. As expected, the closest word to $V_{I}^{k^{\star}}$ “Baby babbling”, however, in the second row (next second of the video) both the women and the baby are emitting sound. As one might have expected that those two sounds were encoded in two different factors, it turns out that they both are encoded on the sounding factor $V_{I}^{k^{\star}}$, as its similarity is maximal with “A baby babbling and a woman talking”, which is coherent with the segmentation. This observation enlightens our previous hypothesis made in section 5.1: the decomposition does not only encode one common concept between the audio and image features, but all the sounding objects, explaining its great performances in the multi-source task.
Figure 5: Multiple source segmentation examples. First row, a baby babbling is correctly segmented. Second row, the sounds of both a baby and a woman are encoded in a single factor showing that our decomposition encodes all sounding objects together.
6 Conclusion
We introduced TACO, a method for training-free sound-prompted segmentation that effectively prompts an open vocabulary image segmentation model such as FC-CLIP with factors extracted through co-factorization of audio-visual deep embedding representations. Our method provides a framework for unsupervised sound-prompted segmentation, enabling the extraction and visualization of shared sounding concepts across audio spectrograms and image regions, which makes it inherently interpretable. TACO significantly outperforms concurrent unsupervised approaches, achieving state-of-the-art results.
Supplementary Material
The supplementary material is organized as follows: the first part presents an analysis of the computations induced by the decomposition method, followed by an explanation of the metrics used and a comparison with other works. The last part consists of qualitative visualization and analysis of segmentation from TACO.
7 Computational overhead
The proposed NMF-based decomposition introduces additional computational costs compared to the baseline (consisting in prompting FC-CLIP with the predicted class from CLAP) as it requires optimization for each batch/sample. Figure 6 shows the performance (mAP) of TACO depending on the number of iterations in the optimization process as well as the associated relative computing time for a complete inference on the ADE SP test set. The performance plateaus after $\sim 1000$ iterations, which takes $\sim 20\%$ additional time compared to the baseline on a single Nvidia A100 80G.
Figure 6: Proportion of additional computing time induce by the decomposition. The decomposition allows an improvement of $\sim 12$ points with only $\sim 20\%$ of additional compute time compared to the baseline.
8 Metrics discussion
[…]
9 Semantic segmentation analysis
Input image TACO prediction Ground Truth segmentation
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Semantic segmentation examples from AVSS.
Figure 7 presents qualitative evidence supporting the hypothesis mentioned in Section 5.2: the semantic segmentation performance is hindered by the class inference process. Specifically, in a scenario with two sound sources, while both sources may be well segmented, each being encoded within the sounding concept through decomposition, the use of a single factor to infer the class leads to both sources being assigned the same class. This significantly penalizes the segmentation score.
10 Qualitative results of single and multiple sources, sound-prompted segmentation
Figure 8 presents randomly sampled segmentation examples from the ADE Sound prompted dataset, as well as their associated activation matrices $U_{I}^{k^{\star}}$, which enlightens the regions of the image that is used to extract the prompt to the segmenter model.
Figure 9 presents the equivalent analysis applied to the multi-sources dataset MS3.
Input image $U_{I}^{k^{\star}}$ TACO prediction Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Sound-prompted segmentation examples from ADE SP.
Input image $U_{I}^{k^{\star}}$ TACO prediction Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Sound prompted-segmentation examples from MS3.
11 Typical failure cases
Input image $U_{I}^{k^{\star}}$ TACO prediction Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Examples of typical failure cases on MS3.
Figure 10 shows typical failure cases of TACO. The first row depicts a case where the decomposition correctly identifies that the concept to segment is a piano (as shown by $U_{I}^{k^{\star}}$), but the segmenter model fails to clearly segment it, resulting in inaccurate segmentation. The second row showcases an example where the decomposition fails to extract a clear concept, as it is unable to encode both audio concepts effectively, yielding partial and inaccurate segmentation. Finally, the third row highlights a special failure case where the decomposition successfully encodes one concept but neglects the second. We hypothesize that this issue arises from limitations in the performance of the audio model (CLAP audio encoder).